OpenStudy (anonymous):

what is the fraction form of 1.2?

OpenStudy (anonymous):

1 1/5

OpenStudy (anonymous):

steps please

OpenStudy (anonymous):

1 and 2/10 which simplifys to 1 and 1/5

OpenStudy (anonymous):

the .2 is equal to 2/10. thats how you get that part

OpenStudy (anonymous):

thank u so much your a great help sr

OpenStudy (anonymous):

think in terms of position value. units . tenths 1 . 2 so 1.2 = 1 +(2/10) (note 1=10/10 as we want everything to the same denominator here 10 ) so 1.2= (10/10) +(2/10) =(12/10) (12/10) =(6/5) (dividing top and bottom by 2) = (5/5) + (1/5) = 1 + (1/5)